OpenStudy (anonymous):

While driving jayant travels 30km with a uniform speed of 40km/h and the next 30km with a uniform speed of 20km/h. Find his average speed.

OpenStudy (anonymous):

find his total time and divide his total distance by it.

OpenStudy (anonymous):

thats it...im confused about it ...pls HELP me.. @Algebraic!

OpenStudy (anonymous):

d= V*t :. t=d/V

OpenStudy (anonymous):

first leg of trip d= 30 km V= 40 km/h t= 30km/(40km /h) = .75 hours

OpenStudy (anonymous):

find the time for the second leg. find the total time. divide the total DISTANCE by the total time.

OpenStudy (anonymous):

time in second leg = 1.5h .....nw i dont knw how to do the division.pls help... @Algebraic!

OpenStudy (anonymous):

@Algebraic!

OpenStudy (anonymous):

pls..........

OpenStudy (anonymous):

add the two times

OpenStudy (anonymous):

add the two distances

OpenStudy (anonymous):

calculate (total distance)/(total time)

OpenStudy (anonymous):

60/1.80????

OpenStudy (anonymous):

srs?

OpenStudy (anonymous):

y?

OpenStudy (anonymous):

what are the times.

OpenStudy (anonymous):

0.75 and 1.5

OpenStudy (anonymous):

add them

OpenStudy (anonymous):

Oops...sry.....2.25